Maricopa's Success scholarship fund receives a gift of $130000. The money is invested in stocks, bonds, and CDs. CDs pay 2.5 % interest, bonds pay 5.5 % interest, and stocks pay 10.5 % interest. Maricopa Success invests $35000 more in bonds than in CDs. If the annual income from the investments is $10600, how much was invested in each account?
Maricopa Success invested $____ in stocks. Maricopa Success invested $____ in bonds. Maricopa Success invested $____ in CDs.

Answer by VFBundy(438)

CDs:
Principal = p
Rate = 0.025
Interest = 0.025p

Bonds:
Principal = p + 35000
Rate = 0.055
Interest = 0.055(p + 35000) = 0.055p + 1925

Stocks:
Principal = 130000 - (p) - (p + 35000) = 95000 - 2p
Rate = 0.105
Interest = 0.105(95000 - 2p) = 9975 - 0.21p

(0.025p) + (0.055p + 1925) + (9975 - 0.21p) = 10600

-0.13p + 11900 = 10600

-0.13p = -1300

p = 10000

CDs:
Principal = p = $10,000

Bonds:
Principal = p + 35000 = 10000 + 35000 = $45,000

Stocks:
Principal = 95000 - 2p = 95000 - 20000 = $75,000